# Crystallization of discrete $N$-particle systems at high temperature

### v1

@@ 1.1. Overview
A similar phenomenon was expected to occur in the discrete counterparts of $\beta$-ensembles. And indeed, it has been confirmed that, for various models, the global limiting distribution is universal in the fixed temperature regime [DF16, BGG17, DS19, Hua21, Mol23, CDM23]. The high temperature regime has also been analyzed recently for three different models [DS19, CDM23, CD25] through the powerful method of moments. The proofs naturally led to various moment-cumulant type formulae, always expressed as weighted generating functions of certain lattice paths, known as Łukasiewicz paths, which generalize well-known moment-cumulant type formulae from classical and free probability.
@@
Although the structure of the combinatorial moment formulas for the limiting measures in the continuous and discrete settings are similar, our simulations in [CD25] suggested that the densities in the discrete setting are qualitatively different; see Fig.˜2. Indeed, we prove that the following crystallization phenomenon of the particles occurs in the high temperature limit: the limiting measures are uniformly supported on disjoint intervals with gaps of unit length. The right endpoints of these intervals turn out to be determined by the zeros of explicit special functions involving Gauss and confluent hypergeometric functions; see Figs.˜3 and 4. We will show that these functions are real-rooted by relating their zeros to the spectra of certain unbounded Jacobi operators. It is worth mentioning that a similar crystallization phenomenon was previously observed in [DS19] for the limit of Jack–Plancherel measures in the high temperature regime, and it was proved in [CDM23] that the locations of the particles in the limit correspond to the zeros of yet another special function, the Bessel function of the first kind. However, the nature of the crystallization in that case was a direct consequence of the scaling of the model, while in the present situation, the fact that the particles are supported on disjoint intervals with unit gaps comes as a surprising consequence of the detailed analysis of the functional equation.
@@ 1.2. LLN for discrete N-particle systems at high temperature and the gamma-quantized R-transform
Theorem 1.1 (Theorem˜4.1 in the text).
@@
Although the combinatorial transform (4) can sometimes be used to show the uniqueness of $\mu$, it is not clear how to use it to study other properties of $\mu$. Some special cases of combinatorial transforms coming from Viennot’s construction [Vie85] are amenable to analysis through the associated orthogonal polynomials, but (4) does not fit into this framework, therefore we needed to find another approach. Our Theorem˜1.1 enables us to analyze the limiting measure using analytic tools, and we apply it to study the high temperature limits of the pure Jack measures to be discussed later. Indeed, Theorem˜1.1 will provide a precise formula for the characteristic function of the limiting measure and will lead to a formula for its density, which displays a remarkable crystallization phenomenon. Additionally, we hope that Theorem˜1.1 will become a key tool to extend the $\gamma$-quantized $R$-transform to measures with heavy tails in future research.
@@ 1.3. Crystallization phenomenon for non-intersecting random walks at high temperature
An important class of examples that can be understood through our $\gamma$-quantized $R$-transform are certain ensembles of discrete $N$-particle random point configurations with log-gas type interaction that depend on the inverse temperature parameter $\beta>0$. One such motivating example derives from the continuous-time, discrete-space, integrable Markov process introduced and studied by Gorin–Shkolnikov [GS15], which has a diffusive limit to the $\beta$-Dyson Brownian motion. This process can be regarded as a one-parameter deformation, non-intersecting $N$-particle version of the Poisson random walk on the integer lattice; see Fig.˜1 for three simulations of this process, with different values of $\theta$.
@@
The construction of [GS15] was based on the Plancherel specialization of the ring of symmetric functions. Later, Huang [Hua21] defined two other variations that result from replacing the Plancherel by the pure alpha and pure beta specializations. They are non-intersecting $N$-particle versions of the geometric and Bernoulli random walks on $\mathbb{Z}$, respectively. The LLN was proved in [Hua21, CD25] for fixed temperature $\theta^{-1}$, general initial conditions, and appropriately chosen sequences of times; this can be appreciated in the left panel of Fig.˜1.
In [CD25], we investigated the high temperature limits of the fixed-time distributions of the previous three families of Markov processes. We proved that for general initial conditions, the answer depends on two ingredients: one is a new operation of probability measures, called the quantized $\gamma$-convolution, explained in [CD25], and secondly, on the limiting measure when the Markov process is started at the trivial initial condition. Moreover, we also proved that the pure Jack measures (to be discussed in the next subsection) are the fixed-time distributions of the Markov processes started at the trivial initial conditions, therefore it remains to find the limiting measures of pure Jack measures. We proved that these limiting measures are uniquely determined by their moments, but solving the moment problem was beyond our reach. Looking at the simulations in the middle and right panels of Fig.˜1, however, we observed that these limiting measures have the following crystallization phenomenon: the large fixed-time distributions of the Markov processes appear to converge to a mixture of uniform distributions supported on the union of countably many intervals. By running a simulation with a larger number of particles and looking at the position of particles at specific times, this phenomenon becomes more visible; see Fig.˜2 below. It seems that the support of the limiting measure has gaps of length $1$, so it is determined by the right endpoints of these intervals. In this paper, we prove that this is indeed true, and we fully describe the crystallization phenomenon for all three cases of pure Jack measures.
@@
Figure 3. Juxtaposition of the simulation (in blue) from Fig.˜2 at time $t\approx\frac{\eta}{\theta}$ and the graph (in green) of the confluent hypergeometric function $_{1}F_{1}(\gamma;z;-\eta)$, where $\eta=1/2$, in the real variable $z$. Theorem˜7.1 predicts that the right endpoints of the intervals on the picture approximate the zeroes of $_{1}F_{1}(\gamma;z;-\eta)$, which are all real.
Figure 4. The same comparison as in Fig.˜3 for $\eta=1$.
@@ Acknowledgements
CC is grateful to the Institute of Mathematics of the Polish Academy of Sciences (IMPAN) for hosting him during his research visits.
@@ 2.3. Jack symmetric functions
is the $Q$-normalization of Jack symmetric functions. In Eq.˜8, we denoted $\lambda^{\prime}$ the conjugate partition to $\lambda$, so that $\lambda_{j}^{\prime}$ is the length of the $j$-th largest column of the Young diagram of $\lambda$.
@@ 3.1. The statement of the LLN
Let $\{\mathbb{P}_{N}\in\mathcal{M}(\mathbb{Y}(N))\}_{N\geq 1}$ be a sequence of probability measures with small tails and let $F_{N,\theta}=F_{N,\theta}(x_{1},\dots,x_{N})$ be the logarithms of their Jack generating functions, defined by Eqs.˜13 and 14. We say that $\{\mathbb{P}_{N}\}_{N\geq 1}$ is HT-appropriate if
@@
A Łukasiewicz path $\Gamma$ of length $\ell$ is a lattice path on $\mathbb{Z}^{2}$ that starts at $(0,0)$, ends at $(\ell,0)$, stays above the $x$-axis, with steps $(1,0)$ (horizontal steps), (1,1)(1,-1) (down steps), and steps $(1,j)$, for some $j\in\mathbb{Z}_{\geq 1}$ (up steps). We denote by $\mathbf{L}(\ell)$ the set of all Łukasiewicz paths of length $\ell$. For example, there are exactly five Łukasiewicz paths of length $3$; see Fig.˜5.
@@
Let $\vec{m}=\mathcal{J}_{\gamma}^{\kappa\mapsto m}(\vec{\kappa})$. From Fig.˜5, the third moment $m_{3}$ is equal to
@@ 3.2. Application to pure Jack measures
where $H_{\theta}(\rho_{1};\rho_{2})$ and $Q_{\lambda}(;\theta)$ are defined by Eqs.˜7 and 8, respectively.
@@
If $\rho$ is a pure alpha, pure beta, or Plancherel specialization, we call $\mathbb{P}_{\rho;N}^{(\theta)}$ a pure Jack measure. Formula (22) allows us to prove the LLN, via Theorem˜3.4, for the three families of pure Jack measures (see [CD25, Sec. 5.2]); in all cases, $\theta=\theta_{N}$ and $\rho=\rho_{N}$ depend on $N$. Let us recall these results.
@@ 3.2.1. Pure alpha Jack measures
The moments of $\mu^{\gamma;c;\eta}_{\operatorname{Alpha}}$ are obtained from these quantized $\gamma$-cumulants and Eq.˜4.
@@ 4. The gamma-quantized R-transform and moment generating function
As a consequence of Theorem˜3.4, we can define, for any $\gamma>0$, the sequence of quantized $\gamma$-cumulants $\{\kappa_{n}^{\gamma}[\mu]\}_{n\geq 1}$ for probability measures $\mu$ that are weak limits, in probability, of atomic measures of the form (16). In fact, the quantities $\kappa_{n}^{\gamma}[\mu]$ are obtained by first computing the moments $m_{n}(\mu)$ and then applying the recursive combinatorial formula from Definition˜3.3. There is certain parallelism to Speicher’s combinatorial approach to Free Probability, where free cumulants of probability measures are obtained from their moments by a recursive formula that involves noncrossing set partitions [Spe94].
@@
Let $\gamma>0$ and assume that $\vec{\kappa}=(\kappa_{1},\kappa_{2},\dots)$ and $\vec{m}=(m_{1},m_{2},\dots)$ are related by the combinatorial identity from Definition˜3.3, i.e. $\vec{m}=\mathcal{J}_{\gamma}^{\kappa\mapsto m}(\vec{\kappa})$. Then there exists a unique $\vec{c}=(c_{1},c_{2},\dots)$ such that the following relations between generating functions hold:
@@
Note that (23) is an equality between power series in $z$, which furnishes formulas between $(\kappa_{1},\kappa_{2},\dots)$ and $(c_{1},c_{2},\dots)$. Likewise, (24) is an equality between power series in $z^{-1}$ that yields formulas between $(c_{1},c_{2},\dots)$ and $(m_{1},m_{2},\dots)$. Put together, these two equalities can be used to express each $m_{\ell}$ in terms of $\kappa_{1},\kappa_{2},\dots$ (and viceversa); Theorem˜4.1 states that this gives an equivalent expression as the one in terms of Łukasiewicz paths from Definition˜3.3.
@@
under certain technical conditions. For example, since $\mathcal{L}\{t^{n}\}(z)=n!\,z^{-n-1}$, for all $n\in\mathbb{Z}_{\geq 0}$ and $\Re z>0$, then $\widetilde{\mathcal{L}}$ coincides with $\mathcal{L}$ for polynomials $f(t)$ in the region $\Re z>0$. In Theorem˜4.1, we decided to use the formal $\widetilde{\mathcal{L}}$ instead of $\mathcal{L}$ to avoid specifying any particular choice of analytic conditions for $f(t)$ and domains for $z$.
@@
Another reason for the interest in Theorem˜4.1 is the (possible) connection with Finite Free Probability, see e.g. [MSS22]. In fact, similar identities were proved in the continuous settings of $\beta$-eigenvalues, see [BGCG22, Thm. 3.11] and $\beta$-singular values, see [Xu23, Thm. 5.8]. It is now known that the identities in those references are precisely the same as in the cumulant-based combinatorial approach to Finite Free Probability started in [AP18] (see also [AGVP23, Cue24]), after the formal identification of the parameter $\gamma\mapsto\textrm{minus the degree of the polynomials}$. Whether our equations (23)–(24) have any relation to Finite Free Probability is an open question.
@@ 4.1. Preliminary lemmas
We need a few lemmas for the proof of Theorem˜4.1. For the first two of them, we denote by $Q^{*}_{\mu}(x_{1},\dots,x_{N};\theta)$, $\mu\in\mathbb{Y}(N)$, the $Q$-normalization of the shifted Jack polynomials, as defined in [OO97] or [OO98, Secs. 2.2–2.4]. In particular, for row partitions $\mu=(k)$, they admit the explicit expression
@@
By Eq.˜26 and the fact that $B_{2m+1}=0$, for all $m\geq 1$, we have
@@
i.e. (34) holds if $f(t)$ is a polynomial and $x>0$. The desired Eq.˜34 is an identity of formal power series in $z^{-1}$, meaning that we want to prove that the coefficients on both sides agree. But such coefficients are just polynomials in $x$ and the coefficients of $f(t)$; since they agree for any $x>0$ and any polynomial $f(t)$, they must also agree for any $x\in\mathbb{R}$ and any formal power series $f(t)$, as needed. ∎
@@
We are finally ready to prove Theorem˜4.1. It should be noted that the main tool will be Theorem˜3.4, which in turn was proved by employing Jack generating functions and Cherednik operators. It would be desirable to have an elementary proof along the same lines as [Cue24, Main Result I]. We leave finding a combinatorial proof as an open problem.
4.2. Proof of Theorem 4.1

### v2

@@ 1.1. Overview
A similar phenomenon was expected to occur in the discrete counterparts of $\beta$-ensembles. And indeed, it has been confirmed that, for various models, the global limiting distribution is universal in the fixed temperature regime [DF16, BGG17, DS19, Hua21, Mol23, CDM23, DGGN25]. The high temperature regime has also been analyzed recently for three different models [DS19, CDM23, CD25] through the powerful method of moments. The proofs naturally led to various moment-cumulant type formulae, always expressed as weighted generating functions of certain lattice paths, known as Łukasiewicz paths, which generalize well-known moment-cumulant type formulae from classical and free probability.
@@
Although the structure of the combinatorial moment formulas for the limiting measures in the continuous and discrete settings are similar, our simulations in [CD25] suggested that the densities in the discrete setting are qualitatively different; see Figure˜2. Indeed, we prove that the following crystallization phenomenon of the particles occurs in the high temperature limit: the limiting measures are uniformly supported on disjoint intervals with gaps of unit length. The right endpoints of these intervals turn out to be determined by the zeros of explicit special functions involving Gauss and confluent hypergeometric functions; see Figures˜3 and 4. We will show that these functions are real-rooted by relating their zeros to the spectra of certain unbounded Jacobi operators. It is worth mentioning that a similar crystallization phenomenon was previously observed in [DS19] for the limit of Jack–Plancherel measures in the high temperature regime, and it was proved in [CDM23] that the locations of the particles in the limit correspond to the zeros of yet another special function, the Bessel function of the first kind. However, the nature of the crystallization in that case was a direct consequence of the scaling of the model, while in the present situation, the fact that the particles are supported on disjoint intervals with unit gaps comes as a surprising consequence of the detailed analysis of the functional equation.
@@ 1.2. LLN for discrete N-particle systems at high temperature and the gamma-quantized R-transform
Theorem 1.1 (Section˜4 in the text).
@@
Although the combinatorial transform (4) can sometimes be used to show the uniqueness of $\mu$, it is not clear how to use it to study other properties of $\mu$. Some special cases of combinatorial transforms coming from Viennot’s construction [Vie85] are amenable to analysis through the associated orthogonal polynomials, but (4) does not fit into this framework, therefore we needed to find another approach. Our Section˜1.2 enables us to analyze the limiting measure using analytic tools, and we apply it to study the high temperature limits of the pure Jack measures to be discussed later. Indeed, Section˜1.2 will provide a precise formula for the characteristic function of the limiting measure and will lead to a formula for its density, which displays a remarkable crystallization phenomenon. Additionally, we hope that Section˜1.2 will become a key tool to extend the $\gamma$-quantized $R$-transform to measures with heavy tails in future research.
@@ 1.3. Crystallization phenomenon for non-intersecting random walks at high temperature
An important class of examples that can be understood through our $\gamma$-quantized $R$-transform are certain ensembles of discrete $N$-particle random point configurations with log-gas type interaction that depend on the inverse temperature parameter $\beta>0$. One such motivating example derives from the continuous-time, discrete-space, integrable Markov process introduced and studied by Gorin–Shkolnikov [GS15], which has a diffusive limit to the $\beta$-Dyson Brownian motion. This process can be regarded as a one-parameter deformation, non-intersecting $N$-particle version of the Poisson random walk on the integer lattice; see Figure˜1 for three simulations of this process, with different values of $\theta$.
@@
The construction of [GS15] was based on the Plancherel specialization of the ring of symmetric functions. Later, Huang [Hua21] defined two other variations that result from replacing the Plancherel by the pure alpha and pure beta specializations. They are non-intersecting $N$-particle versions of the geometric and Bernoulli random walks on $\mathbb{Z}$, respectively. The LLN was proved in [Hua21, CD25] for fixed temperature $\theta^{-1}$, general initial conditions, and appropriately chosen sequences of times; this can be appreciated in the left panel of Figure˜1.
In [CD25], we investigated the high temperature limits of the fixed-time distributions of the previous three families of Markov processes. We proved that for general initial conditions, the answer depends on two ingredients: one is a new operation of probability measures, called the quantized $\gamma$-convolution, explained in [CD25], and secondly, on the limiting measure when the Markov process is started at the trivial initial condition. Moreover, we also proved that the pure Jack measures (to be discussed in the next subsection) are the fixed-time distributions of the Markov processes started at the trivial initial conditions, therefore it remains to find the limiting measures of pure Jack measures. We proved that these limiting measures are uniquely determined by their moments, but solving the moment problem was beyond our reach. Looking at the simulations in the middle and right panels of Figure˜1, however, we observed that these limiting measures have the following crystallization phenomenon: the large fixed-time distributions of the Markov processes appear to converge to a mixture of uniform distributions supported on the union of countably many intervals. By running a simulation with a larger number of particles and looking at the position of particles at specific times, this phenomenon becomes more visible; see Figure˜2 below. It seems that the support of the limiting measure has gaps of length $1$, so it is determined by the right endpoints of these intervals. In this paper, we prove that this is indeed true, and we fully describe the crystallization phenomenon for all three cases of pure Jack measures.
@@
Figure 3. Juxtaposition of the simulation (in blue) from Figure˜2 at time $t\approx\frac{\eta}{\theta}$ and the graph (in green) of the confluent hypergeometric function $_{1}F_{1}(\gamma;z;-\eta)$, where $\eta=1/2$, in the real variable $z$. Section˜7 predicts that the right endpoints of the intervals on the picture approximate the zeroes of $_{1}F_{1}(\gamma;z;-\eta)$, which are all real.
Figure 4. The same comparison as in Figure˜3 for $\eta=1$.
@@ Acknowledgements
CC is grateful to the Institute of Mathematics of the Polish Academy of Sciences (IMPAN) for hosting him during his research visits. MD would like to thank to Sławomir Dinew for stimulating discussions.
@@ 2.3. Jack symmetric functions
is the $Q$-normalization of Jack symmetric functions. In Equation˜8, we denoted $\lambda^{\prime}$ the conjugate partition to $\lambda$, so that $\lambda_{j}^{\prime}$ is the length of the $j$-th largest column of the Young diagram of $\lambda$.
@@ 3.1. The statement of the LLN
Let $\{\mathbb{P}_{N}\in\mathcal{M}(\mathbb{Y}(N))\}_{N\geq 1}$ be a sequence of probability measures with small tails and let $F_{N,\theta}=F_{N,\theta}(x_{1},\dots,x_{N})$ be the logarithms of their Jack generating functions, defined by Equations˜13 and 14. We say that $\{\mathbb{P}_{N}\}_{N\geq 1}$ is HT-appropriate if
@@
A Łukasiewicz path $\Gamma$ of length $\ell$ is a lattice path on $\mathbb{Z}^{2}$ that starts at $(0,0)$, ends at $(\ell,0)$, stays above the $x$-axis, with steps $(1,0)$ (horizontal steps), (1,1)(1,-1) (down steps), and steps $(1,j)$, for some $j\in\mathbb{Z}_{\geq 1}$ (up steps). We denote by $\mathbf{L}(\ell)$ the set of all Łukasiewicz paths of length $\ell$. For example, there are exactly five Łukasiewicz paths of length $3$; see Figure˜5.
@@
Let $\vec{m}=\mathcal{J}_{\gamma}^{\kappa\mapsto m}(\vec{\kappa})$. From Figure˜5, the third moment $m_{3}$ is equal to
@@ 3.2. Application to pure Jack measures
where $H_{\theta}(\rho_{1};\rho_{2})$ and $Q_{\lambda}(;\theta)$ are defined by Equations˜7 and 8, respectively.
@@
If $\rho$ is a pure alpha, pure beta, or Plancherel specialization, we call $\mathbb{P}_{\rho;N}^{(\theta)}$ a pure Jack measure. Formula (22) allows us to prove the LLN, via Section˜3.1, for the three families of pure Jack measures (see [CD25, Sec. 5.2]); in all cases, $\theta=\theta_{N}$ and $\rho=\rho_{N}$ depend on $N$. Let us recall these results.
@@ 3.2.1. Pure alpha Jack measures
The moments of $\mu^{\gamma;c;\eta}_{\operatorname{Alpha}}$ are obtained from these quantized $\gamma$-cumulants and Equation˜4.
@@ 4. The gamma-quantized R-transform and moment generating function
As a consequence of Section˜3.1, we can define, for any $\gamma>0$, the sequence of quantized $\gamma$-cumulants $\{\kappa_{n}^{\gamma}[\mu]\}_{n\geq 1}$ for probability measures $\mu$ that are weak limits, in probability, of atomic measures of the form (16). In fact, the quantities $\kappa_{n}^{\gamma}[\mu]$ are obtained by first computing the moments $m_{n}(\mu)$ and then applying the recursive combinatorial formula from Section˜3.1. There is certain parallelism to Speicher’s combinatorial approach to Free Probability, where free cumulants of probability measures are obtained from their moments by a recursive formula that involves noncrossing set partitions [Spe94].
@@
Let $\gamma>0$ and assume that $\vec{\kappa}=(\kappa_{1},\kappa_{2},\dots)$ and $\vec{m}=(m_{1},m_{2},\dots)$ are related by the combinatorial identity from Section˜3.1, i.e. $\vec{m}=\mathcal{J}_{\gamma}^{\kappa\mapsto m}(\vec{\kappa})$. Then there exists a unique $\vec{c}=(c_{1},c_{2},\dots)$ such that the following relations between generating functions hold:
@@
Note that (23) is an equality between power series in $z$, which furnishes formulas between $(\kappa_{1},\kappa_{2},\dots)$ and $(c_{1},c_{2},\dots)$. Likewise, (24) is an equality between power series in $z^{-1}$ that yields formulas between $(c_{1},c_{2},\dots)$ and $(m_{1},m_{2},\dots)$. Put together, these two equalities can be used to express each $m_{\ell}$ in terms of $\kappa_{1},\kappa_{2},\dots$ (and viceversa); Section˜4 states that this gives an equivalent expression as the one in terms of Łukasiewicz paths from Section˜3.1.
@@
under certain technical conditions. For example, since $\mathcal{L}\{t^{n}\}(z)=n!\,z^{-n-1}$, for all $n\in\mathbb{Z}_{\geq 0}$ and $\Re z>0$, then $\widetilde{\mathcal{L}}$ coincides with $\mathcal{L}$ for polynomials $f(t)$ in the region $\Re z>0$. In Section˜4, we decided to use the formal $\widetilde{\mathcal{L}}$ instead of $\mathcal{L}$ to avoid specifying any particular choice of analytic conditions for $f(t)$ and domains for $z$.
@@
Another reason for the interest in Section˜4 is the (possible) connection with Finite Free Probability, see e.g. [MSS22]. In fact, similar identities were proved in the continuous settings of $\beta$-eigenvalues, see [BGCG22, Thm. 3.11] and $\beta$-singular values, see [Xu23, Thm. 5.8]. It is now known that the identities in those references are precisely the same as in the cumulant-based combinatorial approach to Finite Free Probability started in [AP18] (see also [AGVP23, Cue24]), after the formal identification of the parameter $\gamma\mapsto\textrm{minus the degree of the polynomials}$. Whether our equations (23)–(24) have any relation to Finite Free Probability is an open question.
@@ 4.1. Preliminary lemmas
We need a few lemmas for the proof of Section˜4. For the first two of them, we denote by $Q^{*}_{\mu}(x_{1},\dots,x_{N};\theta)$, $\mu\in\mathbb{Y}(N)$, the $Q$-normalization of the shifted Jack polynomials, as defined in [OO97] or [OO98, Secs. 2.2–2.4]. In particular, for row partitions $\mu=(k)$, they admit the explicit expression
@@
By Equation˜26 and the fact that $B_{2m+1}=0$, for all $m\geq 1$, we have
@@
i.e. (34) holds if $f(t)$ is a polynomial and $x>0$. The desired Equation˜34 is an identity of formal power series in $z^{-1}$, meaning that we want to prove that the coefficients on both sides agree. But such coefficients are just polynomials in $x$ and the coefficients of $f(t)$; since they agree for any $x>0$ and any polynomial $f(t)$, they must also agree for any $x\in\mathbb{R}$ and any formal power series $f(t)$, as needed. ∎
@@
We are finally ready to prove Section˜4. It should be noted that the main tool will be Section˜3.1, which in turn was proved by employing Jack generating functions and Cherednik operators. It would be desirable to have an elementary proof along the same lines as [Cue24, Main Result I]. We leave finding a combinatorial proof as an open problem.
4.2. Proof of Theorem 4